OpenStudy (anonymous):

estimate. then find the sum or difference. 6,292 + 7,318 =

OpenStudy (anonymous):

Estimating to the thousands place 6000 + 7000 = 13000 Or Estimating to the hundreds place 6300 + 7300 = 13600 Solving Exactly 6292+7318= 13610

OpenStudy (anonymous):

13,610